# Data-Driven Discovery of Mobility Periodicity for Understanding Urban Transportation Systems

### v1

@@ 4 Regularity of Ridesharing Trips
Figure 4: Identification of weekly periodicity in ridesharing trip data across different urban areas with the multidimensional sparse autoregression method (sparsity level $\tau=6$). (A) Weekly periodicity of NYC ridesharing pickup trip data across 265 areas from 2019 to 2024. The average periodicity values are also marked. (B) Weekly periodicity of Chicago ridesharing pickup trip data across 77 areas from 2019 to 2024. (C) The coefficient difference between 2024 and 2019, referring to the coefficients of 2024 minus 2019. (D) Annual ridesharing trip counts of NYC and Chicago from 2019 to 2024. In contrast to NYC, the ridesharing trips in Chicago show a poor recovery in the post-pandemic years. Note that the ridesharing trip data of January 2019 is not available from the data portal. (E) Weekly average trip time series of NYC ridesharing, starting from Monday to Sunday. (F) Weekly average trip time series of Chicago ridesharing, starting from Monday to Sunday. While NYC shows the recovery percentage as 68% in 2021, Chicago keeps 44%, roughly as the same level of 2020.
Multidimensional sparse autoregression quantifies human mobility periodicity through coefficients. The resulting non-negative coefficients directly quantify the strength of the positive auto-correlation at certain time lags. A high coefficient at a time lag corresponding to a weekly cycle (e.g., 168 hours for hourly time series data) indicates a strong weekly periodicity.
As we know, the COVID-19 pandemic had a significant disruptive impact on urban mobility regularity and periodicity. To examine how such disruptions affected periodicity, we analyze large-scale ridesharing trip data from 2019 to 2024 in NYC [^2] The NYC ridesharing trip record dataset is available at https://www.nyc.gov/site/tlc/about/tlc-trip-record-data.page. and Chicago [^3] The Chicago ridesharing trip record dataset is available at https://data.cityofchicago.org/.. This analysis aims to demonstrate how the strength of weekly periodicity varied across different urban areas and years, reflecting the impact of the pandemic. In the experiments, we set the sparsity level and the order of sparse autoregression as $\tau=6$ and $d=168$, respectively. On both NYC and Chicago ridesharing datasets, the support set is optimized as $\Omega=\{1,23,24,143,167,168\}$, covering the indices of local auto-correlations, daily cycle, and weekly cycle.
@@
Our analysis so far suggests that the COVID-19 pandemic led to a significant decrease in the periodicity of urban mobility in both NYC and Chicago in 2020. Both cities showed signs of recovery in subsequent years, with periodicity patterns returning to pre-pandemic levels by 2022-2024. However, the pace of recovery in terms of both periodicity and overall trip counts appeared to be faster in NYC compared to Chicago. Furthermore, our method enables policymakers to measure disruptions such as pandemics and design targeted recovery strategies.
@@ 5 Discussion
Understanding the periodicity of human mobility is critical for revealing the underlying structure of individual and collective movement patterns. Periodic patterns in human mobility—ranging from daily commutes to weekly rituals—are manifestations of deeply rooted behavioral rhythms. These patterns are essential components for making informed decisions, developing effective urban policies, and building more resilient and responsive urban systems. In literature, such rhythms have been extensively studied in chronobiology, behavioral sciences, and temporal network theory (Roenneberg et al., 2004; Dijksterhuis et al., 2006; Holme and Saramäki, 2012), where periodicity reveals regularities in decision-making, attention cycles, and social synchronization. Drawing on these insights, we propose a novel method for quantifying and leveraging periodicity in large-scale mobility data. This enables a deeper understanding of urban rhythms and enhances predictive modeling for transportation planning, epidemic spread (Balcan et al., 2009; Belik et al., 2011; Castillo-Chavez et al., 2016), and urban resilience assessments (Xu et al., 2025).
@@
Our results advance people’s understanding of real-world urban systems and show the capability of multidimensional sparse autoregression for quantifying mobility regularity from time series data. Our findings concerning the weekly periodicity—a comparable metric across different time series, spatial areas, and time periods—in NYC and Chicago reveal the disruptive impact of the COVID-19 pandemic on mobility regularity and subsequent recovery trends. The interpretability offered by the multidimensional sparse autoregression is a key advantage, as sparse and non-negative auto-correlations provide actionable insights into the specific temporal lags driving the observed periodicity (Chen et al., 2025b). The method we present allows substantial analysis of time series data that collected from real-world systems and demonstrated periodicity or seasonality. Despite identifying weekly periodicity, the revealed dominant auto-correlations in the support set enable us to introduce proper differencing operations—a simple yet effective solution to address the non-stationarity issue (Chen et al., 2025c).
@@
As the field of human mobility research moves from analytical studies to data-driven experiments designed to directly learn underlying patterns, there is a growing need to formulate the periodicity quantification problem as a sparse identification of dominant positive coefficients in multidimensional time series autoregression. We have developed an interpretable machine learning method for identifying spatial patterns and temporal dynamics that relate to weekly periodicity in urban areas. While this study focuses on metro passenger flow and ridesharing trips in specific cities, a potential limitation is the focus on these modes. Future work could integrate data from other transportation modes, such as public transit, to capture more comprehensive multi-modal behaviors. In terms of potential applications, this interpretable framework provides a robust foundation for discovering and quantifying temporal patterns related to periodicity and seasonality in complicated data and systems beyond human mobility, including climate systems and healthcare time series.
@@ Appendix B Methodology
which can be solved by mixed-integer programming. The support set is therefore optimized $\Omega$ simultaneously.
@@
for any $n\in[N]$ and $\gamma\in[\delta]$. We solve this optimization problem by using quadratic programming.

### v2

@@ 4 Regularity of Ridesharing Trips
Figure 4: Identification of weekly periodicity in ridesharing trip data across different urban areas with the multidimensional sparse autoregression framework at the sparsity level $\tau=6$. (A) Weekly periodicity of NYC ridesharing pickup trip data across 265 areas from 2019 to 2024. The average periodicity values are also marked. (B) Weekly periodicity of Chicago ridesharing pickup trip data across 77 areas from 2019 to 2024. (C) The coefficient difference between 2024 and 2019, referring to the coefficients of 2024 minus 2019. (D) Annual ridesharing trip counts of NYC and Chicago from 2019 to 2024. In contrast to NYC, the ridesharing trips in Chicago show a poor recovery in the post-pandemic years. Note that the ridesharing trip data of January 2019 is not available from the data portal. (E) Weekly average trip time series of NYC ridesharing, starting from Monday to Sunday. (F) Weekly average trip time series of Chicago ridesharing, starting from Monday to Sunday. While NYC shows the recovery percentage as 68% in 2021, Chicago keeps 44%, roughly as the same level of 2020.
Multidimensional sparse autoregression quantifies human mobility periodicity through coefficients at regularity cycles. The resulting non-negative coefficients directly quantify the strength of the positive auto-correlation at certain time lags. A high coefficient at a time lag corresponding to a weekly cycle (e.g., 168 hours for hourly time series data) indicates a strong weekly periodicity. To analyze these periodicity values of different spatial locations over different years or different travel modes, the framework allows the extraction of spatial patterns that related to periodicity, the observations of long-term evolution of mobility regularity, and the discovery of changing behavior across different travel modes. As we know, the COVID-19 pandemic had a significant disruptive impact on urban mobility regularity and periodicity. To examine how such disruptions affected periodicity, we analyze large-scale ridesharing trip data from 2019 to 2024 in NYC [^2] The NYC ridesharing trip record dataset is available at https://www.nyc.gov/site/tlc/about/tlc-trip-record-data.page. and Chicago [^3] The Chicago ridesharing trip record dataset is available at https://data.cityofchicago.org/.. This analysis aims to demonstrate how the strength of weekly periodicity varied across different urban areas and years, reflecting the impact of the pandemic. In the experiments, we set the sparsity level and the order of sparse autoregression as $\tau=6$ and $d=168$, respectively. On both NYC and Chicago ridesharing datasets, the support set is optimized as $\Omega=\{1,23,24,143,167,168\}$, covering the indices of local auto-correlations, daily cycle, and weekly cycle.
@@
Our analysis so far suggests that the COVID-19 pandemic led to a significant decrease in the periodicity of urban mobility in both NYC and Chicago in 2020. Both cities showed signs of recovery in subsequent years, with periodicity patterns returning to pre-pandemic levels by 2022–2024. However, the pace of recovery in terms of both periodicity and overall trip counts appeared to be faster in NYC compared to Chicago. Furthermore, our method enables policymakers to measure disruptions such as pandemics and design targeted recovery strategies.
5 Mobility Periodicity of Multiple Travel Modes
In Manhattan, multi-modal trip records from 2024 reveal distinct weekly periodicity across four popular travel modes, including ridesharing, yellow taxi, subway, and bikesharing (Fig. 5). Through quantifying human mobility regularity with the multidimensional sparse autoregression framework, this analysis aims to demonstrate the differences of weekly periodicity across four travel modes and the underlying spatiotemporal patterns of multi-modal mobility data. As shown in Fig. 5A, total trip volumes decrease in the order of subway (677.5 million), ridesharing (93.1 million), yellow taxi (36.4 million), and bikesharing (28.1 million). The daily time series in Fig. 5B indicates that subway ridership exhibits the most remarkable periodicity, with highly regular weekly cycles sustained throughout the year. In contrast, ridesharing, yellow taxi, and bikesharing display lower periodic structures and greater short-term variability. Fig. 5C further disaggregates bikesharing trips into membership and casual users. Membership trips constitute 23.0 million rides, or approximately 82% of the total, while casual trips account for the remaining 18%. A strong seasonal pattern is evident, with trip counts peaking during the warmer months (May–October) and declining sharply in winter (January, February, December), reflecting the sensitivity of cycling activity to weather conditions.
Figure 5: Trip counts of ridesharing (RS), yellow taxi (YT), subway (SW), and bikesharing (BS) during the whole year of 2024 in Manhattan, one of the most densely populated boroughs of NYC. (A) The total trip counts of ridesharing, yellow taxi, subway, and bikesharing are 93.1 million, 36.4 million, 677.5 million, and 28.1 million, respectively. (B) Trip time series with a daily resolution in Manhattan. (C) Bikesharing trips of membership and causal users across 12 months of 2024. Most bikesharing trip records are related to the membership users.
As shown in Fig. 6A–B, subway and bikesharing trip records are spatially projected onto spatial areas of Manhattan, yielding trip time series at the same spatial resolution as ridesharing and yellow taxi. Treating each travel mode as a variable, the hourly trip time series across spatial areas form a tensor (Fig. 1A). Given the distinct mobility patterns of bikesharing members and casual users, the variable dimension of mobility tensor comprises ridesharing, yellow taxi, subway, bikesharing (member), bikesharing (casual), and bikesharing (all). The resulting dataset is a tensor of size $69\times 6\times 8784$, with the temporal dimension spanning all 8,784 hours of 2024. By using the multidimensional sparse autoregression as described in Fig. 2, the support set is optimized as $\Omega=\{1,23,24,144,167,168\}$ where $k=168$ refers to as a weekly cycle. The auto-correlations at the index $k=168$ allow one to quantify the weekly periodicity of trip time series across different spatial areas and travel modes. Fig. 6C shows that subway trips are more periodic than other travel modes. Although the trips of ridesharing and yellow taxi have the same level of weekly periodicity, ridesharing trips are less periodic in the south areas and more periodic in the north areas of Manhattan than yellow taxi trips. In the bikesharing system, the comparison between membership and causal trips clearly demonstrates more periodic patterns of membership trips. As bikesharing trip volume is less than ridesharing and yellow taxi, the overall weekly periodicity of bikesharing is stronger than ridesharing and yellow taxi.
Figure 6: Periodicity in multi-modal mobility trip data of 2024 in Manhattan with the multidimensional sparse autoregression method at sparsity level $\tau=6$. (A) Subway stations are projected onto 52 areas in Manhattan (i.e., 69 areas in total). (B) Bikesharing stations are projected onto 67 areas in Manhattan. (C) Weekly periodicity of four different travel modes, i.e., ridesharing, yellow taxi, subway, and bikesharing. According to the average periodicity, these travel modes are sorted as subway, bikesharing, ridesharing, and taxi in decreasing order. (D) Daily periodicity in weekday’s multi-modal mobility trip data.
To examine time-varying mobility patterns, we analyze bimonthly trip datasets at an hourly resolution and quantify the weekly periodicity of different travel modes in Manhattan. As reported in Table 2, weekly periodicity is generally lower in the early and late months of 2024 compared to the middle of the year. In particular, mobility regularity in November and December declines remarkably, possibly due to winter weather and holiday-related disruptions. While subway trips exhibit the strongest periodicity overall, they also experience a significant reduction during these months. By contrast, the trips of ridesharing and yellow taxi maintain relatively stable weekly periodicity throughout the year of 2024. Bikesharing usage demonstrates the greatest seasonal sensitivity, with notably weaker periodicity in colder months such as January–April and November–December.
Table 2: Weekly periodicity of bimonthly trip data of four travel modes in Manhattan of 2024. The dataset was divided into 6 bimonthly time phases.
Ridesharing Yellow taxi Subway Bikesharing (member)
January & February 0.142 0.144 0.231 0.131
March & April 0.191 0.198 0.449 0.134
May & June 0.152 0.143 0.289 0.216
July & August 0.183 0.133 0.408 0.219
September & October 0.205 0.195 0.415 0.290
November & December 0.046 0.055 0.136 0.120
In addition, we consider analyzing the daily periodicity of weekday’s trips in Manhattan. The mobility tensor is of size $69\times 6\times 6288$ with the temporal dimension spanning 6,288 hours of 262 weekdays in 2024. The support set in the multidimensional sparse autoregression framework is optimized as $\Omega=\{1,23,24,95,96,120\}$ with sparsity level $\tau=6$ where the index $k=24$ refers to as a daily cycle. Observing the daily periodicity visualized in Fig. 6D, subway trips are more periodic than other travel modes, but the daily periodicity of the north areas is consistently stronger than the south areas of Manhattan. Compared to the consistent weekly periodicity of ridesharing and yellow taxi trips in Fig. 6C, the daily periodicity of yellow taxi trips is stronger than ridesharing as indicated by Fig. 6D. The daily periodicity of causal bikesharing trips is lower than membership trips, but the distinct regularity became a little marginal.
@@ 6 Discussion
Understanding the periodicity of human mobility is critical for revealing the underlying structure of individual and collective movement patterns in urban areas. Periodic patterns in human mobility—ranging from daily commutes to weekly rituals—are manifestations of deeply rooted behavioral rhythms. These patterns are essential components for making informed decisions, developing effective urban policies, and building more resilient and responsive urban systems. In literature, such rhythms have been extensively studied in chronobiology, behavioral sciences, and temporal network theory (Roenneberg et al., 2004; Dijksterhuis et al., 2006; Holme and Saramäki, 2012), where periodicity reveals regularities in decision-making, attention cycles, and social synchronization. Drawing on these insights, we propose a novel method for quantifying and leveraging periodicity in large-scale mobility data. This enables a deeper understanding of urban rhythms and enhances predictive modeling for transportation planning, epidemic spread (Balcan et al., 2009; Belik et al., 2011; Castillo-Chavez et al., 2016), and urban resilience assessments (Xu et al., 2025).
@@
Our results advance people’s understanding of real-world urban systems and show the capability of multidimensional sparse autoregression for quantifying mobility regularity from time series data. Our findings concerning the weekly periodicity—a comparable metric across different time series, spatial areas, and time periods—in NYC and Chicago reveal the disruptive impact of the COVID-19 pandemic on mobility regularity and subsequent recovery trends. In Manhattan, the analysis of daily and weekly periodicity patterns of four different travel modes, including ridesharing, yellow taxi, subway, and bikesharing, demonstrates the rhythm and regularity strength across different spatial areas and bimonthly time phases. The interpretability offered by the multidimensional sparse autoregression framework is a key advantage, as sparse and non-negative auto-correlations provide actionable insights into the specific temporal lags driving the observed periodicity (Chen et al., 2025b). The method we present allows substantial analysis of time series data that collected from real-world systems and demonstrated periodicity or seasonality. Despite identifying weekly periodicity, the revealed dominant auto-correlations in the support set enable us to introduce proper differencing operations—a simple yet effective solution to address the non-stationarity issue (Chen et al., 2025c).
@@
As the field of human mobility research moves from analytical studies to data-driven experiments designed to directly learn underlying patterns, there is a growing need to formulate the periodicity quantification problem as a sparse identification of dominant positive coefficients in multidimensional time series autoregression. To this end, we develop an interpretable machine learning method for identifying spatial patterns and temporal dynamics that relate to weekly periodicity in urban areas. While this study focuses on metro passenger flow, ridesharing trips, and multi-modal travel behaviors in big cities, a potential limitation is how to extend these results to the cities of different scales and complicated transportation modes. Future work could integrate data from other transportation modes, such as public transit, to capture more comprehensive multi-modal behaviors. In terms of potential applications, this interpretable framework provides a robust foundation for discovering and quantifying temporal patterns related to periodicity and seasonality in complicated data and systems beyond human mobility, including climate systems, internet usage data, and web traffic time series.
@@ Appendix B Methodology
which can be solved by mixed-integer optimization. The support set is therefore optimized $\Omega$ simultaneously.
@@
for any $n\in[N]$ and $\gamma\in[\delta]$. We solve this optimization problem by using quadratic optimization.